OpenStudy (anonymous):

A pair of fair dice is rolled once. Suppose that you lose $10 if the dice sum to 4 and win $11 if the dice sum to 3 or 2. How much should you win or lose if any other number turns up in order for the game to be fair?

OpenStudy (dumbcow):

you need the probabilities of sum being 2,3,or4 there are 36 total possible combinations only 1 gives sum of 2 ( 1,1) --> probability = 1/36 two give sum of 3 (1,2) or (2,1) --> probability = 2/36 3 give sum of 4 (1,3) or (3,1) or (2,2) --> probability = 3/36 probability of something else = (36-3-2-1)/36 = 30/36 multiply probabilities by gain or loss, then set equal to 0 to make game fair \[(\frac{1}{36}+\frac{2}{36})*11 + (\frac{3}{36})*(-10) + (\frac{30}{36})*X = 0\] solve for X

OpenStudy (whpalmer4):

There are 3/36 ways to roll a 4, so the expected value is -$10*3/36 There is 1/36 ways to roll a 2, so the expected value is $11*1/36 There are 2/36 ways to roll a 2, so the expected value is $11*2/36 So with the defined outcomes, the expected value is ($11-$10)*3/36 = $1/12 There are 6 outcomes with defined payouts, leaving 36-6=30 others. $x*30/36 = $1/12 Solve for x to get your win or loss for the other numbers.

OpenStudy (whpalmer4):

it appears great minds think alike :-)

OpenStudy (dumbcow):

ahh yes :)

OpenStudy (whpalmer4):

and we got the same answer, too! :-)

OpenStudy (dumbcow):

side note: games in real life tend not to be fair :(

OpenStudy (whpalmer4):

and the scary looking dude urging you to hurry up and roll the dice probably won't let you break out pencil and paper to check the odds :-)

OpenStudy (dumbcow):

agreed

OpenStudy (anonymous):

hi guys would it be a loss then?

OpenStudy (anonymous):

and what would be the answer be the the nearest cents? it cant be in fractions

OpenStudy (whpalmer4):

\[x*\frac{30}{36} = \frac{$1}{12}\]\[x=\frac{$1}{12}*\frac{36}{30}= \]

OpenStudy (whpalmer4):

And yes, you would pay \(x\) if you rolled any number other than 2,3 or 4.

OpenStudy (anonymous):

what is the value of x than?